### v3

# An Explainable AI Approach for Lung Cancer Detection Using Convolutional Neural Networks

Abstract
Early detection of lung cancer is critical to improving survival outcomes. We present a deep learning framework for automated lung cancer screening from chest computed tomography (CT) images with integrated explainability. Using the IQ-OTH/NCCD dataset (1,197 scans across Normal, Benign, and Malignant classes), we evaluate a custom convolutional neural network (CNN) and three fine-tuned transfer learning backbones: DenseNet121, ResNet152, and VGG19. Models are trained with cost-sensitive learning to mitigate class imbalance and are evaluated using accuracy, precision, recall, F1-score, and ROC-AUC. While ResNet152 achieved the highest accuracy (), DenseNet121, although having slightly lower accuracy than the custom CNN, achieved the most balanced precision and recall, which is critical in medical screening. We further apply Shapley Additive Explanations (SHAP) to visualize evidence contributing to predictions, improving clinical transparency. Results indicate that CNN-based approaches augmented with explainability can provide fast, accurate, and interpretable support for lung cancer screening, particularly in resource-limited settings.
1 Introduction
Lung cancer is one of the leading causes of cancer mortality worldwide. Early diagnosis substantially improves survival, yet accurate reading of CT scans is time-consuming and depends on radiologist expertise and availability. In low-resource or rural settings, the shortage of trained radiologists further delays screening and follow-up. Deep learning, particularly convolutional neural networks (CNNs), has demonstrated strong performance in visual recognition tasks and is increasingly used in medical imaging. However, the adoption of AI in clinical workflows requires not only high performance but also transparency and interpretability to support safe decision-making.
This work develops and evaluates an explainable AI (XAI) pipeline for lung cancer detection from CT images. We compare a custom CNN against three widely used transfer learning backbones—DenseNet121, ResNet152, and VGG19—and integrate SHAP to visualize model evidence. Our contributions are:
A complete training and evaluation pipeline on a public, multi-class CT dataset (Normal/Benign/Malignant) with data augmentation and class-imbalance handling.
A comparative study of custom vs. fine-tuned CNNs for lung cancer detection, including DenseNet121, ResNet152, and VGG19.
An explainability layer using SHAP to highlight image regions driving predictions for clinical interpretability.
2 Related Work
Early lung cancer detection in CT imaging has been studied extensively across traditional machine learning and modern deep learning pipelines.
2.1 Traditional Machine Learning Approaches
Prior to the widespread use of CNNs, lung nodule detection relied on handcrafted feature extraction combined with classifiers such as Support Vector Machines (SVM) [valente2016svm], k-Nearest Neighbors, and Random Forests [gonzalez2016rf]. Features often included shape, texture, and intensity statistics, sometimes extracted after segmentation using thresholding or active contour methods. While interpretable, these approaches suffered from limited generalization due to hand-crafted feature design and variability in CT acquisition protocols.
2.2 Deep Learning with CNNs
The introduction of deep CNNs such as AlexNet [krizhevsky2012alexnet], VGG [simonyan2015vgg], ResNet [he2016resnet], and DenseNet [huang2017densenet] transformed medical image analysis by enabling end-to-end feature learning directly from pixel data. Notable works have applied 2D CNNs to slice-based classification [shen2017multicnn] and 3D CNNs for volumetric nodule analysis [zhu20183dcnn]. Transfer learning from ImageNet-pretrained models has been shown to accelerate convergence and improve performance in small medical datasets [tajbakhsh2016transfer]. Public datasets such as LIDC-IDRI [armato2011lidc] and Kaggle Data Science Bowl have served as benchmarks for these methods.
2.3 Advanced Architectures and Attention Mechanisms
More recent works integrate attention modules [wang2017residualattention] or Vision Transformers (ViT) [dosovitskiy2020vit] to enhance global context modeling, which is especially relevant for detecting subtle malignant features across slices. Multi-scale and multi-view CNN frameworks [shen2015multiscale] combine features from different resolutions or anatomical planes to improve robustness.
2.4 Explainable AI in Medical Imaging
Interpretability remains a key requirement for AI adoption in healthcare. Gradient-based saliency methods such as Grad-CAM [selvaraju2017gradcam] and Integrated Gradients [sundararajan2017integrated] provide visual explanations but can be noisy. Model-agnostic methods like LIME [ribeiro2016lime] and SHAP [lundberg2017shap] offer more stable attributions and can be applied to any classifier. Several works [holzinger2019xai] emphasize that explainable AI can help bridge the trust gap between AI predictions and clinician acceptance.
2.5 Our Contribution in Context
Compared to prior studies, our work uniquely combines (i) multiple CNN backbones including a custom architecture and fine-tuned DenseNet, ResNet, and VGG; (ii) cost-sensitive learning to address class imbalance; and (iii) SHAP-based interpretability specifically tailored to multi-class lung cancer classification in CT scans.
3 Materials and Methods
3.1 Dataset
We use the IQ-OTH/NCCD dataset [iqoth_dataset], comprising CT images labeled as Normal (416), Benign (120), and Malignant (561). Figure 1 shows representative samples; Figure 2 illustrates class distribution.
3.2 Preprocessing and Augmentation
All images are resized to . Data augmentation includes random rotations (up to ), center cropping, and random horizontal flips. Images are normalized using ImageNet statistics. To address class imbalance, we employ cost-sensitive learning by weighting the loss to upweight minority classes.
3.3 Model Architectures
Custom CNN.
Our custom CNN stacks convolutional blocks (Conv–BN–ReLU–MaxPool), followed by a flattened representation and fully connected layers with dropout. Table 1 summarizes the key layers and parameter counts.
| Layer | Output Shape | Params |
| Conv(36), BN, ReLU, MaxPool | 168 | |
| Conv(612), BN, ReLU, MaxPool | 660 | |
| Conv(1232), BN, ReLU, MaxPool | 3,488 | |
| Flatten & FC(327685461) + Dropout | 178,951,509 | |
| FC(54613) | 16,386 |
Transfer Learning Backbones.
We fine-tune ImageNet-initialized DenseNet121 [huang2017densenet], ResNet152 [he2016resnet], and VGG19 [simonyan2015vgg] by replacing and training the classifier head (with dropout for regularization). Trainable parameter counts after freezing the feature extractor are listed in Table 2.
| Backbone | Trainable Params |
| DenseNet121 | 144,667 |
| ResNet152 | 525,315 |
| VGG19 | 2,308,711 |
3.4 Training Protocol
Models are trained using Adam optimizer with learning rate , batch size 8, for up to 50 epochs, using stratified splits for train/validation/test. We monitor validation loss for early stopping and select the checkpoint with the best validation performance.
3.5 Evaluation Metrics
We report overall Accuracy, class-averaged Precision, Recall, and F1-score, and plot ROC curves with AUC for each class. Confusion matrices provide per-class error structures.
3.6 Explainability via SHAP
We apply SHAP [lundberg2017shap] to attribute model predictions to input regions. We visualize positive (evidence-for) and negative (evidence-against) contributions for Normal, Benign, and Malignant predictions, supporting clinician validation.
4 Results
4.1 Learning Curves
Figures 3–6 present the training and validation accuracy/loss curves for each model, providing insight into convergence behavior, stability, and potential overfitting.
The Custom CNN (Figure 3) achieved steady improvement in both training and validation accuracy over the first 20 epochs, plateauing thereafter at high performance without significant divergence between training and validation curves. This pattern suggests good generalization and minimal overfitting. The loss curves confirm smooth convergence, with validation loss tracking training loss closely.
DenseNet121 (Figure 4) exhibited a rapid accuracy gain within the first 10 epochs, followed by a slower, steady climb. Validation accuracy remained close to training accuracy throughout, and the loss curves showed consistent downward trends for both, indicating strong feature transfer from ImageNet pretraining and effective fine-tuning.
For ResNet152 (Figure 5), training accuracy reached near-perfect levels early, but validation accuracy was more variable, especially in later epochs. The validation loss curve showed small oscillations rather than a monotonic decrease, possibly reflecting sensitivity to hyperparameters or the high capacity of ResNet152, which can overfit when data is limited.
VGG19 (Figure 6) underperformed relative to the other models. While training accuracy improved steadily, validation accuracy plateaued early and even declined in later epochs, accompanied by unstable validation loss. This divergence points to overfitting and insufficient generalization, likely due to the deeper architecture’s reliance on more data for optimal tuning.
4.2 Quantitative Performance
Table 3 reports overall test set accuracy. While ResNet152 achieved the highest raw accuracy (), Table 4 highlights that DenseNet121 outperformed in balanced metrics, achieving the highest macro-averaged precision (), recall (), and F1-score ().
Custom CNN reached moderate accuracy () but lower recall than DenseNet121, indicating missed minority-class cases. ResNet152’s strong accuracy was paired with less balanced per-class performance, suggesting dominance by majority-class predictions. VGG19 scored lowest across all metrics, aligning with learning curve findings. (Table 4).
| Model | Accuracy (%) |
| Custom CNN | 92.86 |
| DenseNet121 (fine-tuned) | 89.15 |
| ResNet152 (fine-tuned) | 97.30 |
| VGG19 (fine-tuned) | 79.46 |
| Model | Precision (%) | Recall (%) | F1 (%) |
| Custom CNN | 78 | 71 | 73 |
| DenseNet121 (fine-tuned) | 92 | 90 | 91 |
| ResNet152 (fine-tuned) | 86 | 39 | 42 |
| VGG19 (fine-tuned) | 85 | 35 | 38 |
4.3 Confusion Matrices and ROC–AUC
Figure 7 shows confusion matrices for all models. DenseNet121 demonstrated balanced classification across all three categories, with minimal confusion between Benign and Malignant. Custom CNN effectively identified Normal and Malignant cases but struggled more with Benign vs. Malignant separation. ResNet152 classified most cases correctly but occasionally mislabeled Benign as Malignant—potentially problematic in a clinical context. VGG19 had the most off-diagonal errors.
Figure 8 displays per-class ROC curves for Custom CNN and DenseNet121. DenseNet121 achieved especially high AUC values (Malignant ), underscoring its strong sensitivity to cancerous lesions—crucial in screening where false negatives are costly. Custom CNN also achieved high AUCs but trailed slightly, especially for Benign cases.
4.4 Explainability with SHAP
Figure 9 presents SHAP heatmaps for Normal, Benign, and Malignant examples. Pink highlights indicate features contributing positively to predictions; blue indicates features lowering prediction confidence.
For Malignant cases, SHAP emphasized irregular, dense tissue and spiculated margins—consistent with radiological indicators of malignancy. Benign predictions focused on smooth, well-defined nodules, while Normal cases had diffuse low activations with minor vessel emphasis. DenseNet121’s attention consistently aligned with plausible lung regions, while weaker models like VGG19 occasionally fixated on irrelevant peripheral areas, hinting at overfitting to spurious patterns.
5 Discussion
Although ResNet152 reached the highest accuracy, DenseNet121 produced the most balanced precision/recall/F1 and the highest malignant AUC, which can be preferable in clinical screening to minimize false negatives. VGG19 underperformance suggests that deeper VGG stacks are less robust here without extensive tuning or larger training sets. SHAP visualizations provided clinically plausible evidence maps, potentially aiding radiologist review and trust.
Limitations.
The dataset is modest in size and originates from a single source, which may limit generalization across scanners and institutions. The models operate on 2D slices; incorporating 3D context and volumetric aggregation may improve performance. Finally, SHAP is a post-hoc explanation and does not guarantee causal localization.
Future Work.
We plan to (i) extend to multi-center datasets, (ii) evaluate 3D CNNs and vision transformers, (iii) integrate cost-sensitive/focal losses and advanced sampling, (iv) apply federated learning for privacy-preserving training, and (v) compare multiple XAI methods (e.g., Integrated Gradients, Grad-CAM++) with user studies.
6 Conclusion
We presented an explainable CNN-based pipeline for lung cancer detection on CT images. Across custom and transfer learning models, DenseNet121 offered the best balance of clinical metrics, and SHAP visualizations enhanced interpretability. The approach is promising for decision support, especially where radiology expertise is scarce.
Acknowledgements
We would like to express our sincere gratitude to our project supervisor, Er. Bikal Adhikari, for his valuable time, guidance, and support throughout this work. We also thank our project coordinator, Er. Saban Kumar K.C, for his constant encouragement, and Er. Chaitya Shova Shakya, DHOD of the Department of Computer and Electronics, Communication, and Information at Kathford International College of Engineering and Management, for providing the opportunity and environment to carry out this project. We gratefully acknowledge the creators of the IQ-OTH/NCCD dataset for making their data publicly available, which was essential to this research. Finally, we appreciate the support of our friends, families, and all who contributed directly or indirectly to the successful completion of this work.